U (x, y) = 3x3/4 + 3y

a)Calculate MUx and MUy.

b) Do the consumer’s preferences exhibit a diminishing marginal utility for each good?

c) Calculate MRSx,y.

d) Do the consumer’s preferences exhibit a diminishing marginal rate of substitution of x?
Answer #1

Answer

a) Formula:

MUU

Hence MUx = (9/4)x-1/4

MU_y = rac{partial U }{partial y}

MUy = 3

b) Diminishing Marginal utility for x occurs when MUx is decreases as x increases

Here MUx = (9/4)x-1/4

d(MUx)/dx = -(9/16)x-5/4 < 0

Hence, MUx is decreases as x increases. Hence. consumer’s preferences exhibit a diminishing marginal utility for Good x.

Diminishing Marginal utility for y occurs when MUy is decreases as y increases

Here MUy = 3

d(MUx)/dx = 0

Hence, MUy does not decreases as x increases. Hence. consumer’s preferences will not exhibit a diminishing marginal utility for Good y.

Hence, consumer’s preferences will exhibit a diminishing marginal utility for Good x but not for Good y

c)

MRSxy = MUx/MUy = ((9/4)x-1/4)/3 = (3/4)x-1/4

d)

The consumer’s preferences exhibit a diminishing marginal rate of substitution of x If MRSxy decreases as x increases

MRSxy = (3/4)x-1/4

dMRSxy /dx = -(3/16)x-5/4 < 0

Hence MRS decreases as x increases.

So, The consumer’s preferences exhibit a diminishing marginal rate of substitution of x